### v1

@@ Abstract
We introduce GenAI-Powered Inference (GPI), a statistical framework for both causal and predictive inference using unstructured data, including text and images. GPI leverages open-source Generative Artificial Intelligence (GenAI) models—such as large language models and diffusion models—not only to generate unstructured data at scale but also to extract low-dimensional representations that capture their underlying structure. Applying machine learning to these representations, GPI enables estimation of causal and predictive effects while quantifying associated estimation uncertainty. Unlike existing approaches to representation learning, GPI does not require fine-tuning of generative models, making it computationally efficient and broadly accessible. We illustrate the versatility of the GPI framework through three applications: (1) analyzing Chinese social media censorship, (2) estimating predictive effects of candidates’ facial appearance on electoral outcomes, and (3) assessing the persuasiveness of political rhetoric. An open-source software package is available for implementing GPI.
@@ 1 Introduction
This paper illustrates the power of GPI through the following three applications:
@@
GPI addresses this challenge by leveraging the ability of GenAI to generate or replicate unstructured data and to extract low-dimensional vector representations. Machine learning is then applied to these representations to identify a deconfounder, which is a summary of confounding information. Because these representations were used in the actual generative process, GPI eliminates the need to explicitly model the data-generating process or fine-tune the underlying GenAI models. This makes GPI a broadly accessible framework. Our open-source Python package GPI, which is available at https://gpi-pack.github.io/, implements the proposed methodology, further enhancing its applicability.
@@ 1.1 The Proposed Methodological Framework
Before moving on to each application, we provide a brief overview of the proposed methodology. Figure 1 considers two illustrative settings, in which the goal is to estimate the causal effect of treatment $T$ on an outcome $Y$. The treatment may represent some features of unstructured data $\bm{X}$ as in the “text / image as treatment” case (bottom right) or other variables as in the “text / image as confounder” case (top right). In both settings, the treatment-outcome relationship is confounded by the latent features $\bm{U}$ of the unstructured data $\bm{X}$ as well as other observed confounders $\bm{Z}$. A key challenge is to identify these latent confounding features $\bm{U}$ and adjust for them to estimate causal effects.
@@
To achieve its goal, GPI leverages GenAI models to produce unstructured data $\bm{X}$ and extract corresponding internal representations $\bm{R}$. For existing texts or images, the GenAI model is prompted to regenerate the same content, enabling the derivation of consistent internal representations. Using $\bm{R}$ rather than directly modeling the raw data $\bm{X}$ offers three key advantages. First, $\bm{R}$ is a significantly lower-dimensional embedding of the original high-dimensional input. Second, it encodes rich, relevant information distilled from the extensive data used during the GenAI model’s training. Third, by ensuring that the GenAI model is open-source and operates deterministically, GPI facilitates scientific replicability.
We then apply machine learning to the internal representation $\bm{R}$ to estimate a deconfounder $\bm{f}(\bm{R})$, which is a low-dimensional function that approximates the latent confounding structure $\bm{U}$ and is sufficient for causal effect estimation. This step is critical because directly controlling for high-dimensional unstructured data $\bm{X}$ can violate the positivity assumption and lead to biased inference (D’Amour et al.,, 2021). While the deconfounder is not necessarily unique, we formally show that any valid $\bm{f}(\bm{R})$ enables nonparametric identification of causal effects when adjusted for alongside observed confounders $\bm{Z}$. We propose estimating the deconfounder using a deep neural network architecture designed to encode these necessary properties. The appendix provides theoretical justifications tailored to each application setting.
@@ 1.2 Related Work
A growing body of work addresses causal inference with unstructured high-dimensional data, often by either estimating low-dimensional embeddings (e.g., Veitch et al.,, 2020; Pryzant et al.,, 2021; Gui and Veitch,, 2023; Klaassen et al.,, 2024) or modeling the data-generating process using parametric methods such as topic models (e.g., Fong and Grimmer, 2016; Mozer et al., 2020; Roberts et al., 2020; Ahrens et al., 2021; Egami et al., 2022). These approaches typically rely on strong assumptions, such as the bag-of-words model and fixed embedding schemes, and face substantial challenges due to the statistical and computational complexity of high-dimensional unstructured data. This often leads to biased estimates and unreliable inference.
@@ 2.1 Text as Confounder
Because censored and uncensored posts differ systematically, a naive regression of these outcomes on the treatment variable (prior censorship) would yield biased causal estimates. To address this confounding bias, the original analysis adjusts for the content of the focal post, the post date, and each user’s prior censorship history. We adopt the same covariate adjustment strategy within the GPI framework.
@@
To estimate the average treatment effect on the treated (ATT), we apply double machine learning with two-fold cross-fitting (Chernozhukov et al.,, 2018). Standard errors are clustered at the user level, and extreme propensity scores are truncated following the method proposed by Dorn, (2025). Theoretical justifications and further implementation details are provided in Appendix S1.
Figure 2: Estimated effects of prior censorship experience. For each outcome, we estimate the GPI results from LLaMA3 with 8 billion parameters (green) and Gemma3 with 1 billion parameters (red) using the full sample (Full) and the matched sample (Matched). We compare GPI estimates with those of the text matching approach used in the original analysis (purple). The interval bars represent 95% confidence intervals and the standard errors are clustered at the user level.
Figure 2 compares the results from GPI with those from the text matching approach used in the original analysis, which employed coarsened exact matching (Iacus et al.,, 2012) on estimates from the structural topic model (Roberts et al.,, 2016). For comparison, we apply GPI to both the matched sample (628 users, 879 posts) and the full sample (4,155 users; 75,324 posts). In contrast to the original analysis, GPI finds that prior censorship significantly reduces users’ subsequent posting activity, providing evidence of self-censorship. Moreover, GPI reveals a much stronger effect of prior censorship on the likelihood of future censorship. Lastly, GPI’s full-sample estimates closely align with those from the matched sample, but are substantially more efficient, highlighting GPI’s advantages in both robustness and precision. These findings are consistent across both LLaMA 3 and Gemma 3 models.
To further illustrate GPI’s advantage, we assess covariate balance by computing the Pearson’s correlation between the estimated efficient score of GPI and a candidate confounder: the proportion of 60 censorship or self-censorship-related keywords identified by Fu et al., (2013). For text matching, we compute the Pearson’s correlation between this candidate confounder and the residuals from weighted OLS (with weights from coarsened exact matching). If a given method effectively adjusts for the confounding features in focal posts, it should yield a small correlation close to zero with a large $p$-value.
@@
Number of posts $0.004$ $0.073$ $0.002$ $0.081$ $0.032$
$(0.273)$ $(0.094)$ $(0.532)$ $(0.061)$ $(0.398)$
Rate of censorship $-0.001$ $-0.016$ $0.000$ $-0.024$ $0.089$
$(0.783)$ $(0.587)$ $(0.925)$ $(0.489)$ $(0.001)$
Rate of missing posts $0.001$ $0.007$ $-0.000$ $0.042$ $0.062$
$(0.735)$ $(0.826)$ $(0.977)$ $(0.218)$ $(0.063)$
Table 1: Pearson’s correlation (and associated $p$-value in parentheses) between the estimated efficient score of GPI and a candidate confounder (proportion of 60 censorship or self-censorship-related keywords). For text matching, the correlation between the candidate confounder and the residual from the implied weighted OLS is computed. If a method appropriately adjusts for the text features, then the correlation should be small and its $p$-value should be large. “Full” refers to the estimates using the entire sample while “Matched” refers to those based on the matched sample.
As shown in Table 1, GPI yields the correlations that are small in magnitude and have large $p$-values. In particular, the full sample analysis shows that the correlations are precisely estimated to be close to zero with large $p$-values. In contrast, the text matching approach produces a small $p$-value for the rate of censorship variable, despite its lower statistical power due to the limited size of matched sample, suggesting that residual confounding may still remain.
@@ 2.2 Predictive Effects of Image Features
The primary predictors of interest are three facial traits previously identified as socially significant: attractiveness, trustworthiness, and competence (Sutherland et al.,, 2013). To quantify these traits, the authors fine-tuned a pre-trained convolutional neural network using the “One Million Impressions” dataset (Peterson et al.,, 2022). The outcome variable is the standardized number of votes (using $z$-scores) received by each candidate. The original analysis employed an ordinary least squares (OLS) regression of standardized vote counts on the three facial features, along with controls for age, gender, and education.
@@
Each candidate photo is processed using two different versions of Stable Diffusion model (version 1.5 and 2.1) (Rombach et al.,, 2022) to extract its internal representation. We then estimate the deconfounder and outcome model using the same neural network architecture and fine-tuning procedure as in the previous application. Because the original predictors of interest — facial features — are continuous scores, we first discretize them based on 10 bins based on the sample quantiles and estimate the average predict effect at each bin. Theoretical justifications and implementation details are provided in Appendix S2.
Figure 3: Estimated predictive effects of three facial traits (Attractiveness, Dominance, and Trustworthiness) using GPI with Stable Diffusion (SD) version 1.5 (top) and version 2.1 (middle) and OLS (bottom). In each panel, solid and dash lines represent estimates with and without covariate adjustment, respectively. Blue and red shaded areas indicate the corresponding 95% confidence intervals.
Figure 3 compares the GPI estimates to the original OLS results, focusing on the sensitivity of predictive effects to covariate adjustment. Solid and dotted lines represent estimates with and without covariate adjustment, respectively, with blue and red shaded bands indicating the corresponding point-wise 95% confidence intervals. The GPI estimates exhibit greater robustness to the inclusion of covariates. Furthermore, GPI consistently finds that facial dominance is not statistically significantly associated with electoral outcomes, whereas the OLS results vary depending on whether covariates are included. This robustness likely reflects GPI’s ability to account for additional latent features present in candidates’ facial appearance. As before, these findings are consistent across GenAI models.
@@ 2.3 Structural Model of Texts
where $Y_{jj^{\prime}(i)}\in\{0,1,2\}$ denotes the relative persuasiveness of arguments $j$ and $j^{\prime}$ shown to respondent $i$ with 00 indicating that argument $j$ is less persuasive than $j^{\prime}$, $1$ indicating equal persuasiveness, and $2$ indicating argument $j$ is more persuasive, The variable $T_{j}\in\{0,1,\ldots,13\}$ denotes a rhetorical element used in argument $j$, and $\bm{U}_{j}$ represents latent confounding features of that argument. Lastly, $\mu(T_{j},\bm{U}_{j})$ is a strength of the argument $j$, which is a nonparametric function of rhetorical element $T_{j}$ and other confounding features $\bm{U}_{j}$. The original paper uses a special case of this semiparametric formulation, where $\mu(T_{j},\bm{U}_{j})$ is constrained to a linear-additive form and $\bm{U}_{j}$ only captures policy areas and positions.
@@ 3 Concluding Remarks
ATT was then computed using Equation (3), with standard errors derived from the influence function in Equation (S3). To account for within-user correlation, we clustered standard errors at the user level. For consistency with the original study’s text matching approach, we estimated the ATT using both the full sample and the matched sample from the original analysis.
@@
To further assess the robustness of our approach, we examine the correlation between the estimated efficient scores of GPI and a potential confounding variable: the proportion of censorship-related keywords in each post. The literature of semiparametric statistics suggests that the estimated efficient score is mean independent of any observed confounder (e.g., Tsiatis,, 2006). Thus, if the proposed method captures the confounding features in focal posts well, we expect the estimated efficient scores to be uncorrelated with this candidate confoudner.
Specifically, we use a set of 60 keywords identified by Fu et al., (2013): (1) 30 keywords with the highest relative frequency in censored posts compared to uncensored ones, and (2) 30 keywords most frequent among self-censoring users relative to others (as reported in Appendix Tables 1a and 1b of Fu et al., 2013). Then, our candidate confounder is the proportion of these keywords in each post.
For GPI, we directly use the estimated efficient scores from the DML procedure and compute its Pearson’s correlation with the candidate confounder. For text matching, we compute the correlation between the candidate confounder and the residual from its implied weighted OLS (with weights from coarsened exact matching), which is proportional to the efficient score of the weighted OLS model. Finally, we cluster standard errors at the user level to account for within-user dependence.
@@
where $10<\mathbb{P}(\bm{X}_{i}=\bm{x}\mid\bm{Z}_{i}=\bm{z})<$ 1 for all $i=1,\ldots,N$, $\bm{x}\in\mathcal{X}$, and $\bm{z}\in\mathcal{Z}$.
@@
$\bm{P}$$\bm{R}$$\bm{h}_{\bm{\gamma}}(\bm{R})$$\bm{X}$$\bm{U}=\bm{g}_{\bm{U}}(\bm{X})$$T=g_{T}(\bm{X})$$\bm{Z}$$Y$Deep generative model
@@
We reanalyze the dataset originally compiled by Lindholm et al., (2024). In that study, the authors collected facial photographs of Danish political candidates, along with their vote totals and background characteristics — including date of birth, gender, and education — for the 2021 local and 2022 general elections. The initial dataset contained 9,020 candidates. However, the authors excluded individuals whose photos were missing, of low quality, or overlaid with text, as well as those who ran uncontested. This resulted in a final analytic sample of 7,055 candidates. To quantify facial traits, the authors fine-tuned a pre-trained convolutional neural network using the “One Million Impressions” dataset (Peterson et al.,, 2022), generating estimates for three dimensions of perceived facial characteristics: attractiveness, dominance, and trustworthiness.
@@
Unlike in the text-based application, we did not apply a pooling operation to reduce dimensionality. Instead, we flattened the full internal representation (a tensor of shape $16\times 16\times 64$) into a single vector of dimension 16,384, which is substantially smaller than the original image size of $304\times 304\times 3=277,248$.

### v2

@@ Abstract
We introduce GenAI-Powered Inference (GPI), a statistical framework for both causal and predictive inference using unstructured data, including text and images. GPI leverages open-source Generative Artificial Intelligence (GenAI) models—such as large language models and diffusion models—not only to generate unstructured data at scale but also to extract low-dimensional representations that are guaranteed to capture their underlying structure. Applying machine learning to these representations, GPI enables estimation of causal and predictive effects while quantifying associated estimation uncertainty. Unlike existing approaches to representation learning, GPI does not require fine-tuning of generative models, making it computationally efficient and broadly accessible. We illustrate the versatility of the GPI framework through three applications: (1) analyzing Chinese social media censorship, (2) estimating predictive effects of candidates’ facial appearance on electoral outcomes, and (3) assessing the persuasiveness of political rhetoric. An open-source software package is available for implementing GPI.
@@ 1 Introduction
We illustrate the power of GPI through the following three applications:
@@
GPI addresses this challenge by leveraging the ability of GenAI to generate or replicate unstructured data and to extract low-dimensional vector representations that are guaranteed to contain all the relevant information. Machine learning is then applied to these representations to identify a deconfounder, which is a summary of confounding information. Because these representations were used in the actual generative process, GPI eliminates the need to explicitly model the data-generating process or fine-tune the underlying GenAI models. This makes GPI a broadly accessible framework. Our open-source Python package GPI, which is available at https://gpi-pack.github.io/, implements the proposed methodology, further enhancing its applicability.
@@ 1.1 The Proposed Methodological Framework
Before moving on to each application, we provide a brief overview of the proposed methodology. Figure 1 considers two illustrative settings, in which the goal is to estimate the causal effect of treatment $T$ on an outcome $Y$. The treatment may represent some features of unstructured data $\bm{X}$ as in the “text/image as treatment” case (bottom right) or other variables as in the “text/image as confounder” case (top right). In both settings, the treatment-outcome relationship is confounded by the latent features $\bm{U}$ of the unstructured data $\bm{X}$ as well as other observed (structured) confounders $\bm{Z}$. A key challenge is to identify these latent confounding features $\bm{U}$ and adjust for them to estimate causal effects.
@@
To achieve its goal, GPI leverages GenAI models to produce unstructured data $\bm{X}$ and extract corresponding internal representations $\bm{R}$. For existing texts or images, GenAI models are prompted to regenerate the identical content, enabling the derivation of consistent internal representations. Because we configure the GenAI to generate $\bm{X}$ as a deterministic function of $\bm{R}$, the latter is guaranteed to contain all the relevant information used to produce the former. Therefore, GPI eliminates the need to finetune $\bm{R}$, unlike the standard approach that relies upon existing generic embeddings such as BERT (Bidirectional Encoder Representations from Transformers) and openAI’s embedding models.
Using $\bm{R}$ rather than directly modeling the raw data $\bm{X}$ offers several key advantages. First, $\bm{R}$ is a significantly lower-dimensional embedding of the original high-dimensional input. Second, it encodes rich, relevant information distilled from the extensive data used during the GenAI model’s training. Third, by ensuring that the GenAI model is open-source and operates deterministically, GPI facilitates scientific replicability. Finally, it is possible to obtain more statistically efficient estimates by combining different GenAI models.
GPI applies machine learning to the internal representation $\bm{R}$ to estimate a deconfounder $\bm{f}(\bm{R})$, which is a low-dimensional function that approximates the latent confounding structure $\bm{U}$ and is sufficient for causal and predictive effect estimation. This step is critical because directly controlling for high-dimensional unstructured data $\bm{X}$ can violate the positivity assumption and lead to biased inference (D’Amour et al.,, 2021). While the deconfounder is not necessarily unique, we formally show that any valid $\bm{f}(\bm{R})$ enables nonparametric identification of causal and predictive effects when adjusted for alongside observed confounders $\bm{Z}$. We propose estimating the deconfounder using a deep neural network architecture designed to directly encode these necessary properties. The appendix provides theoretical justifications tailored to each application setting.
@@ 1.2 Related Work
A growing body of work addresses causal inference with unstructured high-dimensional data, often by either estimating low-dimensional embeddings (e.g., Veitch et al.,, 2020; Pryzant et al.,, 2021; Gui and Veitch,, 2023; Klaassen et al.,, 2024) or modeling the data-generating process using parametric methods such as topic models (e.g., Fong and Grimmer, 2016; Mozer et al., 2020; Roberts et al., 2020; Ahrens et al., 2021; Egami et al., 2022). These approaches typically rely on strong assumptions, such as the bag-of-words model and fixed embedding schemes, and face substantial challenges due to the statistical and computational complexity of high-dimensional unstructured data. This may lead to biased estimates and unreliable inference.
@@ 2.1 Text as Confounder
Because censored and uncensored posts differ systematically, a naive regression of these outcomes on the treatment variable (prior censorship) is likely to yield biased causal estimates. To address this confounding bias, the original analysis adjusts for the content of the focal post based on a topic model, the post date, and each user’s prior censorship history. We adopt the same covariate adjustment strategy within the GPI framework.
@@
To estimate the average treatment effect on the treated (ATT), we apply double machine learning with two-fold cross-fitting (Chernozhukov et al.,, 2018). Standard errors are clustered at the user level, and extreme propensity scores are truncated following the method proposed by Dorn, (2025). We also construct the optimal combination of the GPI estimates from the two GenAI models based on their estimated influence functions, which can yield the more efficient estimates. Theoretical justifications and further implementation details are provided in Appendix S1.
Figure 2: Estimated effects of prior censorship experience. For each outcome, we estimate the GPI results from LLaMA3 with 8 billion parameters (green) and Gemma3 with 1 billion parameters (red) using the full sample (Full) and the matched sample (Matched). We also present results from an optimally weighted combination of the two models (blue). We compare GPI estimates with those of the text matching approach used in the original analysis (purple). The interval bars represent 95% confidence intervals and the standard errors are clustered at the user level.
Figure 2 compares the results from GPI with those from the text matching approach used in the original analysis, which employed coarsened exact matching (Iacus et al.,, 2012) on estimates from the structural topic model (Roberts et al.,, 2016). For comparison, we apply GPI to both the matched sample (628 users, 879 posts) and the full sample (4,155 users; 75,324 posts).
In contrast to the original analysis, GPI finds that prior censorship significantly reduces users’ subsequent posting activity, providing evidence of self-censorship. Moreover, GPI reveals a much stronger effect of prior censorship on the likelihood of future censorship. Lastly, GPI’s full-sample estimates are substantially more efficient than those based on the matched sample, highlighting GPI’s advantages in statistical efficiency. These findings are consistent across both LLaMA 3 and Gemma 3 models, and their optimal combination on average decreases the standard error by 4.4%.
To further illustrate GPI’s advantage, we assess robustness to a text-based confounder by estimating each method with and without this confounder and calculating the relative absolute bias (i.e., the absolute change in point estimates relative to the original estimates without the text-based confounder). The text-based confounder of interest is measured as the proportion of 60 keywords related to censorship or self-censorship as identified by Fu et al., (2013). If a method effectively adjusts for the confounding features in focal posts, the additional adjustment for the text-based confounder should yield only a small change. We also compute the standard error of the relative absolute bias only for GPI, since such a closed-form analytic formula is not available for text matching.
@@
Number of posts $0.084$ $0.242$ $0.062$ $0.370$ $2.507$
$(0.068)$ $(0.682)$ $(0.067)$ $(3.638)$
Rate of censorship $0.007$ $0.196$ $0.037$ $0.132$ $0.323$
$(0.003)$ $(0.140)$ $(0.033)$ $(0.097)$
Rate of missing posts $0.040$ $0.279$ $0.124$ $0.142$ $0.209$
$(0.049)$ $(0.346)$ $(0.042)$ $(0.212)$
Table 1: Relative absolute bias (the absolute change in point estimates from including a text-based confounder, relative to the original estimates) and its standard error (in parentheses). For the text-based confounder, we use the proportion of 60 keywords related to censorship or self-censorship. Standard errors are calculated only for GPI, as an analytic formula is not available for text matching. “Full” refers to estimates using the entire sample, while “Matched” refers to estimates based on the matched sample.
As shown in Table 1, GPI consistently produces smaller relative absolute bias than text matching. In particular, the full-sample analysis shows that GPI estimates change on average by only about five percent when the text-based confounder is included, whereas text matching estimates change on average by more than 100 percent, suggesting that textual features may not be sufficiently controlled.
@@ 2.2 Predictive Effects of Image Features
The primary predictors of interest are three facial traits previously identified as socially significant: attractiveness, trustworthiness, and competence (Sutherland et al.,, 2013). To quantify these traits, the authors fine-tuned a pre-trained convolutional neural network using the “One Million Impressions” dataset (Peterson et al.,, 2022). The outcome variable is the standardized number of votes (using $z$-scores) received by each candidate. The original analysis employed an ordinary least squares (OLS) regression of standardized vote counts on the three facial features, along with three confounders—age, gender, and education.
@@
Each candidate photo is processed using two different versions of Stable Diffusion model (versions 1.5 and 2.1) (Rombach et al.,, 2022) to extract their internal representations. We then estimate the deconfounder and outcome model using the same neural network architecture and fine-tuning procedure as the one employed in the previous application. Because the original predictors of interest — facial features — are continuous scores, we first discretize them into 10 bins based on the sample quantiles and estimate the average predictive effect within each bin. We optimally combine the GPI estimates obtained from the two GenAI models as done in the previous section. Theoretical justifications and implementation details are provided in Appendix S2.
Figure 3: Estimated predictive effects of three facial traits (Attractiveness, Dominance, and Trustworthiness) using GPI (top) and OLS (bottom). In each panel, solid and dash lines represent estimates with and without covariate adjustment, respectively. Blue and red shaded areas indicate the corresponding 95% confidence intervals.
Figure 3 compares the GPI estimates to the original OLS results, focusing on the sensitivity of predictive effects to additional covariate adjustment (age, gender, and education). Solid and dotted lines represent estimates with and without additional covariate adjustment, respectively, with blue and red shaded bands indicating the corresponding point-wise 95% confidence intervals. The GPI estimates exhibit greater robustness to the inclusion of the covariates. Furthermore, GPI consistently finds that facial dominance is not statistically significantly associated with electoral outcomes, whereas the OLS results vary depending on whether the covariates are included. This robustness likely reflects GPI’s ability to account for additional latent features present in candidates’ facial appearance.
Finally, we emphasize that GPI substantially reduces the dimensionality of input vector when compared to the original image (i.e., colored pixels). In our application, the internal representation of Stable Diffusion (for both versions 1.5 and 2.1) is 16,384, which is almost 100 percent smaller than the dimensionarity of input vector based on the original image, which is 786,432.
@@ 2.3 Structural Model of Texts
where $Y_{jj^{\prime}(i)}\in\{0,1,2\}$ denotes the relative persuasiveness of arguments $j$ and $j^{\prime}$ shown to respondent $i$ with 0 indicating that argument $j$ is less persuasive than $j^{\prime}$, $1$ indicating equal persuasiveness, and $2$ indicating argument $j$ is more persuasive, The variable $T_{j}\in\{0,1,\ldots,13\}$ denotes a rhetorical element used in argument $j$, and $\bm{U}_{j}$ represents latent confounding features of that argument. Lastly, $\mu(T_{j},\bm{U}_{j})$ is a strength of the argument $j$, which is a nonparametric function of rhetorical element $T_{j}$ and other confounding features $\bm{U}_{j}$. The original paper uses a special case of this semiparametric formulation, where $\mu(T_{j},\bm{U}_{j})$ is constrained to a linear-additive form and $\bm{U}_{j}$ only captures policy areas and positions.
@@ 3 Concluding Remarks
When we have internal representations from different LLMs, we can combine the estimates based on them to improve statistical efficiency. The following proposition establishes the asymptotic normality of such combined estimator in the case of the two GenAI models. The extension to more than two GenAI models is straightforward.
Proposition 2 (Asymptotic Normality of the Combined Estimator)
Suppose that we have two different GenAI models that give two different internal representations. Let $\mathcal{D}_{ij}=\{Y_{i},T_{i},\bm{R}_{ij},\bm{Z}_{i}\}$ be the observed data obtained from the $j$th GenAI model for $j=1,2$, where $\bm{R}_{ij}$ is the internal representation obtained from the $j$th GenAI model. Let $\hat{\tau}_{j}$ be the ATT estimators obtained from the $j$th GenAI model and $\psi_{j}(\mathcal{D}_{ij})=\psi_{j}(\mathcal{D}_{ij};\tau,\bm{f}_{j},\mu_{1j},\mu_{0j},\pi_{1j})$ be the influence function for $\hat{\tau}_{j}$ for $j=1,2$, where $\bm{f}_{j}$, $\mu_{tj}$, and $\pi_{1j}$ are the true deconfounder, outcome model, and propensity score model for the $j$th GenAI model, respectively. Under Assumptions 14 for each GenAI model, the combined estimator $\hat{\tau}_{\rm comb}=\omega\hat{\tau}_{1}+(1-\omega)\hat{\tau}_{2}$ for any weight $\omega$ satisfies asymptotic normality:
$\sqrt{N}\frac{(\hat{\tau}_{\rm comb}-\tau)}{\sigma_{\rm comb}}\xrightarrow[]{d}\mathcal{N}(0,1)$
where $\sigma_{\rm comb}^{2}=\mathbb{E}[(\omega\psi_{1}(\mathcal{D}_{i1})+(1-\omega)\psi_{2}(\mathcal{D}_{i2}))^{2}]$. Moreover, the combined estimator achieves the minimum asymptotic variance at the optimal weight
$\omega^{*}=\frac{\mathbb{E}[\psi_{2}(\mathcal{D}_{i2})^{2}]-\mathbb{E}[\psi_{1}(\mathcal{D}_{i1})\psi_{2}(\mathcal{D}_{i2})]}{\mathbb{E}[\psi_{1}(\mathcal{D}_{i1})^{2}]+\mathbb{E}[\psi_{2}(\mathcal{D}_{i2})^{2}]-2\mathbb{E}[\psi_{1}(\mathcal{D}_{i1})\psi_{2}(\mathcal{D}_{i2})]}.$
The proof is given in Appendix S4.2.
@@
ATT was then computed using Equation (3), with standard errors derived from the influence function in Equation (S3). To account for within-user correlation, we clustered standard errors at the user level. For consistency with the original study’s text matching approach, we estimated the ATT using both the full sample and the matched sample from the original analysis. We also consider the optimal combinations of the estimates from the two LLMs using the optimal weights in Theorem 2 so that we can have more efficient results.
@@
GPI (LLaMA3-8B) GPI (Gemma3-1B) GPI (Combined) Text matching
Outcome Full Matched Full Matched Full Matched Matched
Rate of censorship 0.012 0.018 0.011 0.015 0.011 0.017 0.004
(0.000) (0.002) (0.000) (0.002) (0.000) (0.002) (0.001)
Rate of missing posts 0.065 0.100 0.070 0.099 0.067 0.099 0.050
(0.003) (0.022) (0.003) (0.023) (0.003) (0.021) (0.016)
Number after event -16.200 -14.700 -14.400 -4.330 -15.600 -7.690 11.500
(2.600) (21.900) (2.690) (21.300) (2.560) (21.100) (41.200)
Table S2: Estimated Average Treatment Effects for the Treated (ATT) using GPI with LLaMA3 (8B), Gemma3 (1B), their optimal combination, and text matching (Roberts et al.,, 2020). Standard errors (reported in parentheses) are clustered at the user level. The table presents results for both the full sample and the matched sample.
@@
S1.3 Examining the Robustness to the Inclusion of Additional Covariates
To further assess the robustness of our approach, we examine how much the point estimate changes when we additionally adjust for the textual features that might act as a confounder. Specifically, we estimate both the GPI and text matching methods with and without directly controlling the text features, and calculate the relative absolute bias (i.e., the absolute change in point estimates divided by the absolute value of the original point estimates). For GPI, we can further calculate the standard error of the relative absolute bias using the delta method, as shown below.
Formally, let $\tau_{\mathrm{long}}$ and $\tau_{\mathrm{short}}$ be the parameter of interest obtained from GPI with and without controlling the text features, respectively, and let $C_{i}$ be the text-based confounding feature for post $i$. Then, the relative absolute bias is defined as
$R=\frac{|\tau_{\mathrm{long}}-\tau_{\mathrm{short}}|}{|\tau_{\mathrm{short}}|}.$
Let $\psi_{\mathrm{long}}$ and $\psi_{\mathrm{short}}$ be the influence functions of $\tau_{\mathrm{long}}$ and $\tau_{\mathrm{short}}$, respectively. As they are influence function, we have
$\sqrt{N}\begin{pmatrix}\hat{\tau}_{\mathrm{long}}-\tau_{\mathrm{long}}\\ \hat{\tau}_{\mathrm{short}}-\tau_{\mathrm{short}}\end{pmatrix}=\frac{1}{\sqrt{N}}\sum_{i=1}^{n}\begin{pmatrix}\psi_{\mathrm{long}}(\mathcal{D}_{i},C_{i};\tau_{\mathrm{long}},\bm{f}_{\mathrm{long}},\mu_{1,\mathrm{long}},\mu_{0,\mathrm{long}},\pi_{\mathrm{long}})\\ \psi_{\mathrm{short}}(\mathcal{D}_{i};\tau_{\mathrm{short}},\bm{f}_{\mathrm{short}},\mu_{1,\mathrm{short}},\mu_{0,\mathrm{short}},\pi_{\mathrm{short}})\end{pmatrix}+o_{P}(1)$
where $\bm{f}_{\mathrm{long}}$, $\mu_{t,\mathrm{long}}$, and $\pi_{\mathrm{long}}$ are the nuisance functions estimated when controlling the text features, and $\bm{f}_{\mathrm{short}}$, $\mu_{t,\mathrm{short}}$, and $\pi_{\mathrm{short}}$ are those estimated without controlling the text features. By the multivariate central limit theorem, we have
$\sqrt{N}\begin{pmatrix}\hat{\tau}_{\mathrm{long}}-\tau_{\mathrm{long}}\\ \hat{\tau}_{\mathrm{short}}-\tau_{\mathrm{short}}\end{pmatrix}\xrightarrow{d}\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}\sigma_{\mathrm{long}}^{2}&\sigma_{\mathrm{long,short}}\\ \sigma_{\mathrm{long,short}}&\sigma_{\mathrm{short}}^{2}\end{pmatrix}\right)$
where $\sigma_{\mathrm{long}}^{2}=\mathbb{E}[\psi_{\mathrm{long}}^{2}]$, $\sigma_{\mathrm{short}}^{2}=\mathbb{E}[\psi_{\mathrm{short}}^{2}]$, and $\sigma_{\mathrm{long,short}}=\mathbb{E}[\psi_{\mathrm{long}}\psi_{\mathrm{short}}]$. Then, under the assumption that $\tau_{\mathrm{short}}\neq 0$ and $\tau_{\mathrm{long}}-\tau_{\mathrm{short}}\neq 0$, by the delta method, we have
$\sqrt{n}(\hat{R}-R)\xrightarrow{d}\mathcal{N}(0,V)$
where
$V=\nabla g^{\top}\begin{pmatrix}\sigma_{\mathrm{long}}^{2}&\sigma_{\mathrm{long,short}}\\ \sigma_{\mathrm{long,short}}&\sigma_{\mathrm{short}}^{2}\end{pmatrix}\nabla g$
with g(x,y)=|xy|/|y|g(x,y)=|x-y|/|y| and $\nabla g=(\partial g/\partial x,\partial g/\partial y)^{\top}$ evaluated at $(\tau_{\mathrm{long}},\tau_{\mathrm{short}})$.
To calculate the relative absolute bias, we use a set of 60 keywords identified by Fu et al., (2013): (1) 30 keywords with the highest relative frequency in censored posts compared to uncensored ones, and (2) 30 keywords most frequent among self-censoring users relative to others (as reported in Appendix Tables 1a and 1b of Fu et al., 2013). Then, our candidate confounder is the proportion of these keywords in each post.
@@
where $0<\mathbb{P}(\bm{X}_{i}=\bm{x}\mid\bm{Z}_{i}=\bm{z})<1$ for all $i=1,\ldots,N$, $\bm{x}\in\mathcal{X}$, and $\bm{z}\in\mathcal{Z}$.
@@
$\bm{P}$$\bm{R}$$\bm{h}_{\bm{\gamma}}(\bm{R})$$\bm{X}$$\bm{U}=\bm{g}_{\bm{U}}(\bm{X})$$T=g_{T}(\bm{X})$$\bm{Z}$YYDeep generative model
@@
We reanalyze the dataset originally compiled by Lindholm et al., (2024). In that study, the authors collected facial photographs of Danish political candidates, along with their vote totals and background characteristics — including date of birth, gender, and education — for the 2021 local and 2022 general elections. The initial dataset contained 9,020 candidates. However, the authors excluded individuals whose photos were missing, of low quality, or overlaid with text, as well as those who ran uncontested. In addition, while most of the photos are square, we found that 237 photos have a different aspect ratio. To ensure consistency for the downstream analysis, we excluded these photos from the dataset. This resulted in a final analytic sample of 6,868 candidates. To quantify facial traits, the authors fine-tuned a pre-trained convolutional neural network using the “One Million Impressions” dataset (Peterson et al.,, 2022), generating estimates for three dimensions of perceived facial characteristics: attractiveness, dominance, and trustworthiness.
@@
Unlike in the text-based application, we did not apply a pooling operation to reduce dimensionality. Instead, we flattened the full internal representation (a tensor of shape $16\times 16\times 64$) into a single vector of dimension 16,384, which is substantially smaller than the original image size of $512\times 512\times 3=786,432$.